OpenStudy (anonymous):

What's the sum of the roots of sin x (2sin x - 1) = 2sin x - 1? Can someone check my work and help me find out where I'm making mistakes? sin x (2sin x - 1) = 2sin x - 1 2 sin²x - sin x = 2sin x - 1 2 sin² x -3sin x + 1 = 0 Let sin x be A 2A² - 3A + 1 = 0 (2A - 1)(A - 1) = 0 So, sin x = 1/2 or sin x = 1

OpenStudy (anonymous):

P.S. The interval is \[0 \le x \le pi\]

OpenStudy (anonymous):

So I suppose if sin x = 1/2, then x must be pi/6 or (6/5)pi And if sin x = 1, then x must be pi/2

OpenStudy (anonymous):

Anyway, my book says the sum is 3pi/2

OpenStudy (kinggeorge):

If \(\sin(x)=1/2\), then \[x={\pi \over 6} \quad \text{or}\quad x={5\pi \over 6}\]

OpenStudy (kinggeorge):

Not \[6\pi \over 5\]

OpenStudy (anonymous):

Oh, I'm so silly. Thanks!

OpenStudy (kinggeorge):

You're welcome.